# Neurosymbolic Grounding for Compositional World Models

### v1

@@ A.4 Downstream Evaluation Setup
Following [36], we use a greedy planner that chooses the action that minimizes the Hungarian distance between the current and the goal state. These actions are applied $t-1$ times over a trajectory of length $t$, with the output from the world model at the ($d-1$)-th step becoming the state for step $d$ in the trajectory. Due to this compounding nature, we see an increased divergence from the ground truth as we get deeper into the trajectory. At each step $d$ in the trajectory, the accuracy of the world model is evaluated as the L1 error of the difference between the current ground truth and predicted states in the form of their $xy$-coordinates. These $xy$-coordinates are initialized for each object to the origin and updated with every action taken by the corresponding rule. For example, after one step, if the ground truth moves an object to the east, but the planner chooses to move the same object to the west, then the distance between the two states would be 2. We run these experiments for the 500 trajectories of length $t=32$ in our test dataset and average the scores at each trajectory depth. We showcase results in Figure 4. Our model (Cosmos) shows the most consistency and least deviation from the goal state in all datasets, which suggests that neurosymbolic grounding helps improve the downstream efficacy of world models.
A.5 Qualitative Results

### v2

@@ A.5 Downstream Evaluation Setup
Following Veerapaneni et al. (2020), we use a greedy planner that chooses the action that minimizes the Hungarian distance between the current and the goal state. These actions are applied $t-1$ times over a trajectory of length $t$, with the output from the world model at the ($d-1$)-th step becoming the state for step $d$ in the trajectory. Due to this compounding nature, we see an increased divergence from the ground truth as we get deeper into the trajectory. At each step $d$ in the trajectory, the accuracy of the world model is evaluated as the L1 error of the difference between the current ground truth and predicted states in the form of their $xy$-coordinates. These $xy$-coordinates are initialized for each object to the origin and updated with every action taken by the corresponding rule. For example, after one step, if the ground truth moves an object to the east, but the planner chooses to move the same object to the west, then the distance between the two states would be 2. We run these experiments for the 500 trajectories of length $t=32$ in our test dataset and average the scores at each trajectory depth. We showcase results in Figure 4. Our model (Cosmos) shows the most consistency and least deviation from the goal state in all datasets, which suggests that neurosymbolic grounding helps improve the downstream efficacy of world models.
A.6 Dataset Comparison
The focus of our paper is to demonstrate the first neuro-symbolic framework leveraging foundation models for compositional object-oriented world modelling, and we evaluate on the same benchmarks as existing work (Kipf et al., 2019; Goyal et al., 2021; Zhao et al., 2022). We chose our evaluation domain based on four properties: (1) object-oriented state and action space, (2) history-invariant dynamics, (3) action conditioned (plannable) dynamics, and (4) ease of generating new configurations (to evaluate entity and relational composition). We curate the following list of domains from related work to explain what properties are missing for each dataset.
Dataset and Relevant Works Object Oriented State and Action Space History Independent Action Conditioned (Plannable) Configurable 2D Block Pushing (Kipf et al., 2019; Goyal et al., 2021) (Zhao et al., 2022; Ke et al., 2021) Physion (Wu et al., 2023) Phyre (Wu et al., 2023) Clevrer (Wu et al., 2023) 3DObj (Wu et al., 2023) MineRL (Hafner et al., 2023) Atari (Pong/Space invaders/Freeway) (Goyal et al., 2021) Minigrid/BabyAI Mao et al. (2022)
A.7 Symbolic Ablation of Cosmos
AlignedNPS serves as a “fully neural” ablation to demonstrate the effectiveness of our model. In this section, we detail another “fully symbolic” ablation of our model to demonstrate the need for a neurosymbolic approach. Specifically, we maintain the algorithm presented in 1 but modify the transition model to use the symbolic embedding to predict the next state. Specifically, line 9 changes to:
$S_{p}$ $\leftarrow S_{p}+\textsc{MlpBank}[r](\textbf{concat}(\Lambda_{p},~{}\Lambda_{c},~{}\vec{R}_{r}))$
The results, detailed in Table 3, indicate that the “symbols-only” model significantly underperforms compared to Cosmos. We believe this is because the symbolic embedding is constructed by concatenating symbolic attributes, and the rule module is not aware of this structure. This causes the MLP to overfit to the attribute compositions seen at train time. Cosmos sidesteps this issue by using the symbolic embedding in the key-query attention module to select the relevant rule module, while allowing the real vector to learn local features useful for modeling action-conditioned transitions.
Dataset Only Symbols (MSE $\downarrow$) Only Neural / AlignedNPS (MSE $\downarrow$) COSMOS (MSE $\downarrow$) 3 Object RC - Sticky 1.36E-02 1.14E-02 4.23E-03 3 Object RC - Team 1.39E-02 1.24E-02 4.60E-03 3 Object EC 1.21E-02 3.51E-03 7.66E-04
Table 3: Evaluation results on the 2D block pushing domain for ablations of Cosmos.
A.8 Qualitative Results